# Toroidal Cartesian Products Where One Factor is $3$-Connected

### v1

@@ 3. The genus of $K_{4}\boxempty P_{3}$
Suppose $f_{3}=6$, $f_{4}=6$, and $f_{5}=2$. By Lemma 3.1, each $K_{4}$ fiber contains exactly two triangles. Note that, each pentagon borders a pair of edges in either $E_{1}$ or $E_{2}$. Thus, all six quadrilaterals must be mixed quadrilaterals in order to make sure each edge in $E_{1}\cup E_{2}$ borders exactly two faces. Each mixed quadrilateral borders exactly one edge in $K^{2}$, each pentagon borders either one or two edges in $K^{2}$, and there are two triangles bordering a total of six edges in $K^{2}$. As a result, a minimum of $14$ edges in $K^{2}$ are bordered by faces, which forces some edge in $K^{2}$ to border at least three faces, providing a contradiction.
@@
Suppose $f_{3}=6$, $f_{4}=7$, and $f_{6}=1$. By Lemma 3.1, each $K_{4}$ fiber contains exactly two triangles. As discussed before the proof, there are three types of hexagons that can exist in an embedding of $G$. We now break this case into sub-cases, depending on what type of hexagon is used.
@@
We now describe how to embed $G$ on $S_{2}$. First, embed each $K_{4}$ fiber on a separate $2$-sphere as the $1$-skeleton of a tetrahedron. We use $F_{i}$ for $i=1,\ldots,4$ to denote the four triangles in one of these embeddings. First, delete the interiors of $F_{1}$ and $F_{2}$ from the $K^{1}$ embedding and from the $K^{2}$ embedding. Insert a tube (homeomorphic to $\mathbb{S}^{1}\times[0,1]$), between the $F_{j}$ face deleted from $K^{1}$ and the $F_{j}$ face deleted from $K^{2}$ for $j=1,2$. The four edges of $E_{1}$ can be embedded on these two tubes. Now, delete the triangular faces $F_{3}$ and $F_{4}$ coming from $K^{2}$’s embedding on a $2$-sphere and also delete $F_{3}$ and $F_{4}$ from $K^{3}$’s embedding on a $2$-sphere. Insert two tubes in a similar fashion to embed the edges of $E_{2}$, completing the embedding of $G$. Since this embedding involves three disjoint $2$-spheres joined by four tubes, the resulting genus of the surface is two. ∎
4. Restricting toroidal $G\boxempty H$ via vertex connectivity.
@@
We proceed by contrapositive. Suppose $\kappa(G)\geq 5$. Since $H$ is a simple connected graph with $|V(H)|\geq 2$, we see that $H$ contains a $P_{2}$ subgraph. Note that, $|V(G\boxempty P_{2})|=2|V(G)|$ and $|E(G\boxempty P_{2})|=2|E(G)|+|V(G)|$. Since $G$ is simple, the girth of $G\boxempty P_{2}$ is at least three. By Proposition 2.1, we have that
@@
Since $G$ is $5$-connected, for all $v\in V(G)$, we have $deg(v)\geq 5$. Since $G$ is simple, every edge is incident to two distinct vertices. Thus, $\frac{5}{2}|V(G)|\leq|E(G)|$, and so, we can conclude that
$\gamma(G\boxempty P_{2})\geq 1+\frac{1}{3}(\frac{5}{2}|V(G)|)-\frac{5}{6}|V(G)|=1.$
Equality holds if and only if $G$ is $5$-regular and $G\boxempty P_{2}$ admits a triangular emdedding. However, any edge in a $P_{2}$ fiber is not contained in any $3$-cycle of $G\boxempty P_{2}$, and so, a triangular embedding of $G\boxempty P_{2}$ does not exist. Therefore, $\gamma(G\boxempty H)\geq\gamma(G\boxempty P_{2})>1.$
@@
The work of Tutte [18] shows that any simple, $3$-connected graph contains a wheel graph $W_{n}$ as a minor, for some $n\geq 3$; see Figure 2 for a diagram of $W_{5}$. At the same time, $W_{m}$ is a minor of $W_{n}$, whenever $m\leq n$. It is easy to see that one can first delete an appropriate number of the spoke edges of $W_{n}$ that connect to the central vertex, and then, contract some of the edges along the $n$-cycle of $W_{n}$ to obtain a $W_{m}$ minor. Thus, $W_{3}=K_{4}$ is a minor of any simple, $3$-connected graph, as needed. ∎
@@
We proceed by contrapositive. Suppose $H$ is a (simple, connected, non-trivial) graph where $H\neq P_{2}$. Then $H$ contains a $P_{3}$ subgraph. By Lemma 4.2, $G$ has a $K_{4}$ minor. Then by Theorem 3.4 and Corollary 2.3, we have that $\gamma(G\boxempty H)\geq\gamma(K_{4}\boxempty P_{3})=2$. ∎
Combined, Theorem 4.3 and Proposition 4.1 give rise to the following questions:
Question 4.4.
Does there exist an infinite set of $3$-connected (44-connected) graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal? If so, can we classify such $G_{i}$?
5. Constructing toroidal $G\boxempty P_{2}$, where $\kappa(G)\in\{3,4\}$.
In this section, we address the first part of Question 4.4 and provide a conjecture for the second part of Question 4.4. This first requires us to introduce a new class of graphs. A graph $G$ is outer-cylindrical if it embeds on the $2$-sphere such that there are two distinct faces whose boundaries collectively contain all the vertices of $G$. Since the class of outer-cylindrical graphs is closed under taking minors, the Graph Minor Theorem of Robertson and Seymour [16] implies that there is a finite list of minor-minimal obstructions for this class of graphs, which were then determined by Archdeacon–Bonnington–Dean–Hartsfield–Scott [2]. Following their notation, we will use the abbreviation OC to stand for outer-cylindrical and NOC to stand for non-outer-cylindrical. The following proposition motivates our interest in OC graphs.
Proposition 5.1.
Let $G$ be an OC-graph. Then $\gamma(G\boxempty P_{2})\leq 1$.
@@
We first construct disjoint OC-embeddings of the two $G$ fibers (call them $G_{1}$ and $G_{2}$) of $G\boxempty P_{2}$ in the $2$-sphere $\mathbb{R}\cup\{\infty\}$. Take an OC-embedding of $G$ as $G_{1}$, and reflect $G_{1}$ over a line that does not intersect $G_{1}$ to build the OC-embedding of $G_{2}$. All the vertices of each $G_{i}$ for $i=1,2$ are contained in two faces of this embedding, and one of these faces, which we call the outer face, is common between $G_{1}$ and $G_{2}$. For each vertex $v$ in $G_{1}$ on this outer face, we can embed on edge on the $2$-sphere connecting $v$ to the corresponding vertex in $G_{2}$. We can then cut out the interiors of the “inner” faces for $G_{1}$ and $G_{2}$ and add a tube between them carrying the remaining edges of $G\boxempty P_{2}$ to build an embedding of $G\boxempty P_{2}$ on the torus.∎
@@
Figure 1. The inner and outer faces for the $G_{1}$ fiber are depicted in green and the inner and outer faces for the $G_{2}$ fiber are depicted in black. The edges contained in the $P_{2}$ fibers that connect to vertices on the two outer faces are depicted in blue. The edges contained in the $P_{2}$ fibers that connect to vertices on the inner faces are not depicted.
It is easy to see that there exist infinite classes of $3$-connected graphs that are OC. For this, we recall the definition of a wheel of order n for $n\geq 3$, denoted $W_{n}$, as the simple connected graph with $n+1$ vertices and $2n$ edges with the following properties:
  1. (1)
    $V(W_{n})=\{c,b_{0},\ldots,b_{n-1}\}$,
  2. (2)
    $c$ is adjacent to each $b_{i}$ for $i=0,\ldots n-1$,
  3. (3)
    $b_{i}$ and $b_{j}$ are adjacent if and only if $|i-j|=1\mod n$.
@@
Figure 2. The standard planar embedding of $W_{5}$ with vertices labeled.
Proposition 5.2.
For $n\geq 3$, each wheel $W_{n}$ is OC. Consequently, there exists an infinite set of OC, $3$-connected graph $\{W_{n}\}_{n=3}^{\infty}$ such that each $W_{n}\boxempty P_{2}$ is toroidal.
Proof.
For $n\geq 3$, consider the standard planar embedding of $W_{n}$ as given in Figure 2. Then the outer region is a face bound by the $n$-cycle $b_{0}b_{1},\ldots,b_{n-1}$. The triangle $cb_{0}b_{1}$ bounds $c$, and so, these two faces contain every vertex, which implies that this embedding is OC. The work of Tutte [18] shows that each wheel is simple and $3$-connected. By Proposition 5.1, each $W_{n}\boxempty P_{2}$ is toroidal. ∎
We can also construct $4$-connected graphs that are OC. For $n\geq 4$, define $C^{2}_{n}$ as the graph obtained from the $n$-cycle $C_{n}$ by adding an edge between every pair of vertices that are a distance two apart in $C_{n}$. See Figure 3 for a standard planar embedding of $C^{2}_{6}$.
[…]

### v3

@@ 3 The genus of $K_{4}\boxempty P_{3}$
Suppose $f_{3}=6$, $f_{4}=6$, and $f_{5}=2$. By Lemma 1, each $K_{4}$ fiber contains exactly two triangles. Note that, each pentagon borders a pair of edges in either $E_{1}$ or $E_{2}$. Thus, all six quadrilaterals must be mixed quadrilaterals in order to make sure each edge in $E_{1}\cup E_{2}$ borders exactly two faces. Each mixed quadrilateral borders exactly one edge in $K^{2}$, each pentagon borders either one or two edges in $K^{2}$, and there are two triangles bordering a total of six edges in $K^{2}$. As a result, a minimum of $14$ edges in $K^{2}$ are bordered by faces, which forces some edge in $K^{2}$ to border at least three faces, providing a contradiction.
@@
Suppose $f_{3}=6$, $f_{4}=7$, and $f_{6}=1$. By Lemma 1, each $K_{4}$ fiber contains exactly two triangles. As discussed before the proof, there are four types of hexagons that can exist in an embedding of $G$. We now break this case into sub-cases, depending on what type of hexagon is used.
@@
We now describe how to embed $G$ on $S_{2}$. First, embed each $K_{4}$ fiber on a separate $2$-sphere as the $1$-skeleton of a tetrahedron. We use $F_{i}$ for $i=1,\ldots,4$ to denote the four triangles in one of these embeddings. First, delete the interiors of $F_{1}$ and $F_{2}$ from the $K^{1}$ embedding and from the $K^{2}$ embedding. Insert a tube (homeomorphic to $\mathbb{S}^{1}\times[0,1]$), between the $F_{j}$ face deleted from $K^{1}$ and the $F_{j}$ face deleted from $K^{2}$ for $j=1,2$. The four edges of $E_{1}$ can be embedded on these two tubes. Now, delete the triangular faces $F_{3}$ and $F_{4}$ coming from the $K^{2}$ embedding on a $2$-sphere and also delete $F_{3}$ and $F_{4}$ from the $K^{3}$ embedding on a $2$-sphere. Insert two tubes in a similar fashion to embed the edges of $E_{2}$, completing the embedding of $G$. Since this embedding involves three disjoint $2$-spheres joined by four tubes, the resulting genus of the surface is two. ∎
4 Restricting toroidal $G\boxempty H$ via vertex connectivity.
@@
We proceed by contrapositive. Suppose $\kappa(G)\geq 5$. Since $H$ is a simple connected graph with $|V(H)|\geq 2$, we see that $H$ contains a $P_{2}$ subgraph. Note that, $|V(G\boxempty P_{2})|=2|V(G)|$ and $|E(G\boxempty P_{2})|=2|E(G)|+|V(G)|$. Since $G$ is simple, the girth of $G\boxempty P_{2}$ is at least three. By Proposition 1, we have that
@@
Since $G$ is $5$-connected, for all $v\in V(G)$, we have $deg(v)\geq 5$. By the Handshaking Lemma, $\sum deg(v)=2|E(G)|$. Thus, $\frac{5}{2}|V(G)|\leq|E(G)|$, and so, we can conclude that
$\gamma(G\boxempty P_{2})\geq 1+\frac{1}{3}\left(\frac{5}{2}|V(G)|\right)-\frac{5}{6}|V(G)|=1.$
Equality holds if and only if $G$ is $5$-regular and $G\boxempty P_{2}$ admits a triangular emdedding. However, any edge in a $P_{2}$ fiber is not contained in any $3$-cycle of $G\boxempty P_{2}$, and so, a triangular embedding of $G\boxempty P_{2}$ does not exist. Corollary 1 implies that $\gamma(G\boxempty H)\geq\gamma(G\boxempty P_{2})>1.$
@@
The work of Tutte Tu1961 shows that any simple $3$-connected graph contains a wheel graph $W_{n}$ as a minor, for some $n\geq 3$. Whenever $m\leq n$, it is easy to see that $W_{m}$ is a minor of $W_{n}$: one can first delete an appropriate number of the spoke edges of $W_{n}$ that connect to the central vertex, and then, contract some of the edges along the $n$-cycle of $W_{n}$ to obtain a $W_{m}$ minor. Thus, $W_{3}=K_{4}$ is a minor of any simple $3$-connected graph, as needed. ∎
@@
We proceed by contrapositive. Suppose $H$ is a (simple, connected, non-trivial) graph where $H\neq P_{2}$. Then $H$ contains a $P_{3}$ subgraph. By Lemma 3, we know that $G$ has a $K_{4}$ minor. Then by Theorem 3.1 and Corollary 1, we have that $\gamma(G\boxempty H)\geq\gamma(K_{4}\boxempty P_{3})=2$. ∎
Combined, Proposition 5 and Proposition 4 give rise to the following questions: Does there exist an infinite set of 33-connected (44-connected) graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal? If so, can we classify such $G_{i}$?
5 Classifying toroidal $G\boxempty P_{2}$, where $\kappa(G)\in\{3,4\}$.
In this section, we fully address both of questions from above. Recall that a graph $G$ is outer-cylindrical if it embeds in the plane or the $2$-sphere such that there are two distinct faces whose boundaries collectively contain all the vertices of $G$. Since the class of outer-cylindrical graphs is closed under taking minors, the Graph Minor Theorem of Robertson and Seymour RoSe2004 implies that there is a finite list of minor-minimal obstructions for this class of graphs, which were then determined by Archdeacon–Bonnington–Dean–Hartsfield–Scott ArBo2001 . The following proposition motivates our interest in outer-cylindrical graphs.
Proposition 6.
If $G$ is an outer-cylindrical graph, then $\gamma(G\boxempty P_{2})\leq 1$.
@@
We first construct disjoint outer-cylindrical embeddings of the two $G$ fibers (call them $G_{1}$ and $G_{2}$) of $G\boxempty P_{2}$ in the $2$-sphere. Take an outer-cylindrical embedding of $G$ as $G_{1}$, and reflect $G_{1}$ over a line that does not intersect $G_{1}$ to build the outer-cylindrical embedding of $G_{2}$. All the vertices of each $G_{i}$ for $i=1,2$ are contained in two faces of this embedding, and one of these faces, which we call the outer face, is common between $G_{1}$ and $G_{2}$. For each vertex $v$ in $G_{1}$ on this outer face, we can embed on edge on the $2$-sphere connecting $v$ to the corresponding vertex in $G_{2}$. We can then cut out the interiors of the “inner” faces for $G_{1}$ and $G_{2}$ and add a tube between them carrying the remaining edges of $G\boxempty P_{2}$ to build an embedding of $G\boxempty P_{2}$ on the torus. See Figure 1 for a schematic of this embedding. ∎
@@
Figure 1: The inner and outer faces for the $G_{1}$ fiber are depicted in green and the inner and outer faces for the $G_{2}$ fiber are depicted in black. The edges contained in the $P_{2}$ fibers that connect to vertices on the two outer faces are depicted in blue. The edges contained in the $P_{2}$ fibers that connect to vertices on the inner faces are not depicted.
In fact, a converse to Proposition 6 also holds.
Proposition 7.
Suppose $G$ is a (connected) graph. If $\gamma(G\boxempty P_{2})\leq 1$, then $G$ is an outer-cylindrical graph.
Proof.
Let $G$ be a (connected) graph such that $\gamma(G\boxempty P_{2})\leq 1$. First, if $\gamma(G\boxempty P_{2})=0$, then the classification of planar Cartesian products of graphs given by Behzad and Mahmoodian BeMa1969 implies that $G$ is outerplanar, and so, $G$ is outer-cylindrical. Moving forward, we assume $\gamma(G\boxempty P_{2})=1$, which implies that $G$ is not outerplanar. Furthermore, there exists a minimal genus embedding of $G\boxempty P_{2}$ on the torus, $\mathbb{T}$, which must be a $2$-cell embedding. We work with this fixed embedding for the remainder of this proof. Let $G^{+}$ and $G^{-}$ be the two $G$-fiber subgraphs of $G\boxempty P_{2}$. Consider the induced embedding of $G^{+}$ on $\mathbb{T}$ and the collection of regions $\mathbb{T}\setminus G^{+}=\{R_{1},\ldots,R_{k}\}$. There are two possible cases to consider depending on the topology of these regions.
Case 1: Suppose each region $R_{i}$, for $1\leq i\leq k$, is a $2$-cell region. Since $G^{-}$ is connected, it must be embedded in the interior of some $2$-cell region $R_{i}$. Since $G^{+}$ is not outerplanar, there exists some vertex $v^{+}\in V(G^{+})$ that is not on the boundary walk $W$ of the region $R_{i}$. However, there exists an edge $e=v^{+}v^{-}$ in a $P_{2}$ fiber connecting $v^{+}$ to the corresponding vertex $v^{-}\in V(G^{-})$, and clearly, $W$ provides an obstruction for $e$ to exist in such an embedding. This gives a contradiction.
Case 2: Suppose there exists some region, say $R_{j}$, that is a non-$2$-cell region. First, $R_{j}$ must be the unique non-$2$-cell region in this collection, since otherwise, $G$ would not be connected. Then either (i) $R_{j}$ is homeomorphic to a once-punctured torus or (ii) $R_{j}$ is homeomorphic to an open annulus. Sub-case (i) implies that $G^{-}$ must be embedded in $R_{j}$, since otherwise, $\gamma(G\boxempty P_{2})=0$. This also implies that $G^{+}$ is embedded on a closed disk in $\mathbb{T}$ whose boundary is some walk in $G^{+}$. However, we can now apply the same argument used in Case 1 to see that there is a vertex on the interior of this closed disk that must connect via an edge to a vertex in the complement of this closed disk, but such an edge would not exist in this embedding, giving a contradiction for this sub-case. So, suppose $R_{j}$ is homeomorphic to an annulus. Then $G^{-}$ must be embedded in $R_{j}$, since otherwise, $\gamma(G\boxempty P_{2})=0$. Let $W_{1}$ and $W_{2}$ be the two boundary closed walks of $R_{j}$. Similar to the argument in Case 1, if there exists $v^{+}\in V(G^{+})$ that is not on $W_{1}\cup W_{2}$, then $W_{1}\cup W_{2}$ forbids the existence of the edge $e=v^{+}v^{-}$, giving a contradiction. Thus every vertex of $G^{+}$ must lie on $W_{1}\cup W_{2}$, and so, $G$ is outer-cylindrical, as needed. ∎
Collecting our major results provides a proof of our classification theorem that was stated in the introduction.
Proof of Theorem 1.1.
Let $G$ be a $3$-connected graph. For the forwards direction, Propositions 5 and 7 collectively give the desired conclusions. For the backwards direction, Proposition 6 gives the desired conclusion. ∎
To conclude this section, we provide one set of examples of Cartesian products of graphs that are toroidal.
Example 1.
For $n\geq 3$, define $C^{2}_{2n}$ as the graph obtained from the $2n$-cycle $C_{2n}$ by adding an edge between every pair of vertices that are a distance two apart in $C_{2n}$. See Figure 2 for a standard planar embedding of $C^{2}_{6}$, which also shows this graph is outer-cylindrical. For $n\geq 3$, $C^{2}_{2n}$ is known to be $4$-connected; see Ma1980 . Thus, $\{C^{2}_{2n}\boxempty P_{2}\}_{n=3}^{\infty}$ provides an infinite class of toroidal Cartesian products of graphs where one factor is $4$-connected. Furthermore, by Propositions 4 and 5, we see that $G=C^{2}_{2n}$ realizes the maximal vertex connectivity for toroidal $G\boxempty P_{2}$.